# NovelQA: A Benchmark for Long-Range Novel Question Answering

### v1

@@ 3.1 Dataset Description
Table 1: Distribution of Question Types in NovelQA: This table provides a breakdown of questions across different complexity categories (Multi-hop, Single-hop, Detailed) and aspect categories (Times, Meaning, Span, Setting, Relation, Character, Plot).
3.2 Data Collection and Annotation
Procedure Overview The annotation process is performed by a group of skilled annotators, all of whom are holding or pursuing a degree in English Literature and have read the target novels carefully, and they are aware of the usage of the data. The annotation procedure consists of two phases: (1) Template-based phase: The annotators are required to fill entities into 19 templates (see Appendix A.1) that we design to be related to multi-hop or detailed information. This phase entails half of the data, mainly contributing to the multi-hop ones. (2) Free-formed phrase: To ensure the diversity of question expression and align the questions to the natural distribution, our second half of the data is annotated without a template, namely, the annotators contribute to any difficult questions that they come up with freely.
Template Design The first annotation phase relies on a question template, which requires the annotator to fill in the entities from the novel to form valid questions. To design templates, we carried out sufficient pre-tests on GPT-4-128K and Claude2-100K to analyze their possible weaknesses in long-input QA and novel knowledge memorization. Our pre-test shows that they usually fail to tackle information spanning over multiple chapters, as well as lack attention to details that have no contribution to the main theme. Meanwhile, we also refer to around fifteen books on novel and narration theories (e.g. Forster, 1927; Tobias, 2012; Schmidt, 2012; McKee, 2005) to ensure our template covers more aspects that a novel can discuss (e.g., character, setting, theme). Aggregating all the information above, the template consisting of the 19 kinds of questions is formed.
Annotation Guideline While creating the QA tuples, our annotators are instructed to follow several principles below: (1) The annotators are either senior-grades English language & literature college students, or students with high English test scores. This ensures the annotators’ proofreading ability. (2) The annotators are required to read through the GPT-4 responses on example questions, ensuring them to be familiar with the task goal. (3) The annotators should choose novels above 50K tokens and read through the books they choose before annotation. (4) Evidence of each answer should be provided for validation purposes. (5) The annotation reward, of \$1.11 per tuple, is explicitly informed in the guidelines. Data of high quality are rewarded by \$1.39 per tuple. As an average annotator can write 5 to 6 pieces of data at full speed according to our observation, the \$5.56 to \$8.34 hourly wage is above the local legal minimum wage of \$2.78/hour.
Quality Control The created data is manually double-checked by authors. Data with factual errors or questions that are too simple are either modified or eliminated. Only 79.4% of the data collected are preserved to form the final dataset of 2305 QA tuples. Meanwhile, the inter-annotator agreement test, carrying on the multichoice versioned QA annotation on randomly selected 100 cases, shows a score of 94.66% in Cohen’s Kappa, indicating a high agreement among annotators.
Distractions for Multi-choice Setting We use GPT-4 to generate three distracting options for each question and its golden answer and randomly permute the four answers.
3.3 Advantages
4 Experiments
Max Length Multi- choice Gene- rative
GPT-4 128K 71.80 46.88
Claude 2.1 200K 66.84 46.04
InternLM2 -Chat-7b 200K 43.51 30.90
InternLM2 -Chat-20b 200K 49.18 32.37
Table 2: Evaluation of Long-Context LLMs on NovelQA. This table presents the performance of four long-context LLMs, including both commercial models (GPT-4 and Claude 2.1) and open-source, locally deployed models (InternLM2-Chat-7b/20b). Accuracy percentages are reported under two testing scenarios: multichoice and generative. The Max Length column denotes the maximum token length of each model.
chara mean plot relat settg span times avg
mh 57.81 61.76 52.46 45.30 56.52 18.18 21.23 32.83
sh 66.12 56.56 69.33 21.43 57.23 - - 63.93
dtl 52.63 12.87 61.68 37.50 58.06 - - 37.58
avg 62.04 32.40 65.93 41.72 57.38 18.18 21.23 46.88
(a) GPT-4
mh 48.94 64.29 58.18 40.00 82.61 18.52 17.10 30.34
sh 72.41 55.24 67.96 23.08 59.60 - - 65.47
dtl 55.22 12.43 66.30 27.59 55.77 - - 37.65
avg 65.90 31.61 66.60 35.61 61.06 18.52 17.10 46.04
(b) Claude 2.1
mh 23.81 38.24 42.62 24.35 39.13 15.15 21.32 24.62
sh 35.80 28.10 42.18 14.29 32.10 - - 36.42
dtl 26.67 9.18 55.14 29.03 34.43 - - 27.02
avg 32.02 18.52 44.77 24.38 33.33 15.15 21.32 30.90
(c) InternLM2-Chat-7b
mh 32.84 44.12 29.03 37.61 25.00 15.15 26.81 29.29
sh 42.97 34.17 43.22 21.42 36.90 - - 40.57
dtl 30.93 7.00 48.21 25.00 29.03 - - 24.65
avg 38.50 19.77 42.66 33.74 33.86 15.15 26.81 33.07
(d) InternLM2-Chat-20b
Table 3: Model Performance by Question Type in Generative Setting: This table details the accuracy scores of four models across different question types. Question types include details (dtl), multi-hop (mh), single-hop (sh), and character (chara), meaning (mean), plot, relation (relat), setting (settg), others, and an average score (avg) for each category, with ’-’ indicating the absence of data for a category.
4.1 Implementations
Settings To thoroughly test the abilities of these LLMs, we employ two evaluation settings: a generative setting where models directly generate short answers, and a multichoice setting with four provided options.
Prompts We use uniform prompts for all LLMs, with a start part, novel content, questions, choices in the multichoice setting, and end part. The prompt structure is shown in Appendix Table 9.
Truncation Due to input length limitations, we truncated the novel content from the end to the front, to meet with the max input length, while keeping questions and other prompts complete.
Evaluating Generative Results: Following the findings in Wang et al. (2023a) which highlight GPT-4’s proficiency in assessing the accuracy of short machine-generated answers, we employ GPT-4 (gpt-4-0125-preview) for the evaluation of generative responses in our study, which is also applied in other long-range benchmark studies (An et al., 2023; Li et al., 2023a). We further conducted a human evaluation on 800 pieces of generative outputs and carried an inter-evaluator agreement test between two human evaluators and the GPT-4 evaluator. The result of 89.25% in Cohen’s Kappa indicates a high agreement towards the GPT-4 evaluating results (details in Appendix B.1).
Commercial LLMs: The APIs of commercial LLMs utilized are gpt-4-0125-preview [^3] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo and anthropic.claude-v2:1444https://aws.amazon.com/cn/bedrock/claude/.
Open-source LLMs: Running long-context LLMs on extremely long inputs, such as 200k tokens, is a challenge due to the immense GPU memory required, for example, it takes roughly 2.5T memory to calculate one attention matrix for a 7B model with a 200K-token input, while our local device is a 4 $\times$ 80G A100. To address this, we utilize the LMDeploy555 https://github.com/InternLM/lmdeploy (based on Dynamic NTK emozilla (2023)) for memory reduction, which is only compatible with several LLMs. Therefore, we choose InternLM2-Chat-7b-200K and InternLM2-Chat-20b-200K for our experiments. Attempts to run models like Yarn-Mistral-7b-128K (Peng et al., 2023), chatglm3-6b-128k (Zeng et al., 2022), XVERSE-13B-256K (xverse, 2023) and Yi-6B-200K (01-ai, 2023) on inputs exceeding 100k tokens are unsuccessful. The challenges range from GPU memory overflow due to incompatibility with LMDeploy, to models imitating the input content, or continuing to generate following the inputs without answering the questions. However, reducing input lengths to under 50k tokens can normalize most outputs, following the given instructions. This situation underscores the current limitations in technical support for processing and generating content based on extremely long contexts.
Other Details Given the cost of running long-context APIs, we request that each model respond to all questions for a given book in a single session. To ensure fair comparisons, local-deployed LLMs also answer all questions for a book at once. We set ‘temperature = 0’ to eliminate randomness and keep other hyper-parameters default.
4.2 Main Results
We present the main results in Table 2. Even the highest scores (71.80% and 46.88% for GPT-4 in generative and multichoice settings, respectively) suggest there is considerable room for improvement in long-context understanding. This is especially true in the generative setting where understanding and recall over long contexts are more challenging. Additionally, commercial models (GPT-4 and Claude 2.1) outperform open-source models (InternLM2-Chat-7b and 20b) in this benchmark. All models show a drop in performance in the generative setting compared to the multichoice setting. This indicates the increased challenge in generating a correct answer from scratch, as opposed to selecting from provided options. We have also observed three typical errors, hallucination, overlooking, and miscounting, a detailed analysis is conducted in Appendix B.3.
4.3 Results by the Question Type
An in-depth analysis of model performance across question types reveals nuanced insights into their comprehension abilities in both generative and multichoice settings, detailed in Tables 3 and Appendix 10, respectively. This analysis not only highlights the models’ weaknesses across different formats but also illuminates the challenges in narrative comprehension, contributing to both NLP and computational literary research.
1. Meaning Questions: The struggle with meaning questions indicates a challenge in grasping abstract concepts and locating entities or sentences through interpretations within the text, which requires an advanced level of semantic understanding and inference.
2. Relation Questions: Difficulty with relation questions points to a gap in the models’ ability to identify and interpret the dynamic and often nuanced relationships between characters, events, or concepts, which are crucial for a holistic understanding of narratives.
3. Span and Times Questions: The lower performance on span and times questions suggests a limitation in tracking temporal sequences and spatial extents within the narrative, reflecting challenges in maintaining and applying contextual information over long stretches of text.
Figure 3: Analysis of Accuracy in Generative Setting by Token Indexes: This figure illustrates the accuracy, plotted against the token indexes of relevant evidences in the novels. The x-axis, reflecting token indexes, is folded on the right due to the long-tail distribution.
@@ 4.3 Results by the Question Type
Figure 4: Analysis of Accuracy in Generative Setting by Relative Token Positions: This figure illustrates the accuracy, plotted against the percentage position of each question’s relevant evidence within the novel.
4.4 Results by the Position
Absolute Position Analysis: In the generative setting, as depicted in Figure 3, all evaluated models show improved performance on questions where the necessary evidence is located before the 100k token mark. This trend highlights a challenge for LLMs in accessing and processing information beyond this threshold, suggesting a diminished capacity to handle very long inputs. The multichoice setting, detailed in Appendix B.2, follows a similar pattern, reinforcing the importance of evidence position in model performance.
Relative Position Analysis: By normalizing the evidence positions within the entire novel, we aim to understand if the proportional location of evidence influences model accuracy. This analysis, shown in Figure 4 for the generative setting and in the Appendix B.2 for the multichoice setting, indicates that models maintain relatively consistent performance across various relative positions, suggesting that long-context LLMs’ effectiveness is not significantly affected by the evidence’s relative position within the standardized length of novels.
4.5 Analysis
@@
Correct- ness Rele- vance Suffi- ciency Avg.
GPT-4 29.47 38.05 27.67 31.73
Cluade 2.1 23.51 29.08 22.27 24.95
InternLM2- Chat-7b 2.05 3.88 1.90 2.61
IntnerLM2- Chat-20b 6.36 12.51 7.50 8.79
Table 5: Evidence recalling results of four LLMs.
@@
Close-book Question Answering We employ a Close-book QA scenario to assess the extent to which models rely on the content of novels versus using their internal knowledge to answer questions. In this approach, the models are not given access to the text of the novels and must rely solely on their pre-existing knowledge to provide answers. Given that our selected novels are well-known and representative of their genres, it’s inevitable that LLMs have encountered their texts during training and retained some of their content. The results of this evaluation are presented in Table 4.
@@
Models like GPT-4 and Claude 2.1 achieve notable scores in the Close-book setting (60.94% and 51.77% in multichoice, 34.30% and 22.36% in generative, respectively) indicating that they have internalized significant portions of the novels’ content during training. This internal knowledge allows them to perform reasonably well even without direct access to the text. The difference in performance between the Close-book and standard settings underscores the challenges in long-context understanding. While models can retain and recall information from well-known texts, their ability to comprehend and use such information to answer questions accurately diminishes in the absence of the text. This suggests that long-context understanding, as measured by the main results, might still be more challenging than it appears, as models benefit from having the text directly available.
Evidences Recall We evaluate the ability to recall evidence, prompting the four models above to answer the questions in NovelQA again, with printing the supporting evidence simultaneously. We then prompt GPT-4 with the generated evidence alongside the annotated evidence to obtain its evaluation on the quality of retrieved evidence pieces. The evaluating matrix consists of the following three dimensions: correctness refers to whether the retrieved evidence is the same as the annotated evidence or with a similar correct meaning; relevance indicates whether the evidence is consistent with the answer; sufficiency, whether the retrieved pieces of evidence are enough to support the answer. Each dimension is scored between 0 and 100 and an average score is further obtained through calculating the algorithmic mean on these three dimensions. Prompts involved in this evaluation procedure are presented in Appendix A.4.
5 Conclusion
We introduced NovelQA, a long-range benchmark designed to assess the long-context comprehension abilities of LLMs. Utilizing representative English novels, NovelQA presents LLMs with the challenge of navigating complex, real-world texts. Our evaluations reveal that both commercial and open-source models face difficulties with detailed understanding, multi-hop reasoning, and accurately retrieving specific information from lengthy contexts, especially for lengths are 100,000. Moreover, operating LLMs on inputs exceeding 200,000 tokens faces technical challenges, notably in terms of memory requirements and associated costs. NovelQA’s contributions extend to advancing research in both NLP and computational literary studies.
Limitations
Appendix
Appendix A Data
Model Human Evaluator A Human Evaluator B Avg.
GPT-4 91.97 95.84 93.91
Claude 2.1 88.04 90.00 89.02
InternLM-Chat-7b 91.88 86.53 84.88
InternLM-Chat-20b 85.68 84.08 87.25
Avg. 89.39 89.11 89.25
A.1 Question Templates
Aspect Template Answer Format
Has the word or phrase <a word or phrase> appeared in the novel? If so, how many times does it appear in the text?
Yes/No + number
Has the object <an object> appeared in the novel? If so, how many times does it appear in the text?
Yes/No + number
times
Has the place <a place> appeared in the novel? If so, how many times does it appear in the text?
Yes/No + number
Has the sentence <a sentence> appeared in the novel? If so, how many times does this sentence appear in the text?
Yes/No + number
Has the character <a character> appeared in the novel? If so, how many times does he/she/they appear in the text?
Yes/No + number
Has the plot <a plot> happened in the novel? If so, how many times does it happen in the text?
Yes/No + number
How many times has <a character> done <a doing-verb phrase> in the novel?
Yes/No + number
How many times have <a character> and <a character> <do something> together in the novel?
Yes/No + number
How many times have <a character> and <a character> <met each other (or appear together)> in the novel?
Yes/No + number
How many times have <a character> and <a character> <communicate or have verbal conflicts with each other> in the novel?
Yes/No + number
mean
Explain the meaning or implication of the symbol or metaphor <a symbol or a metaphor> in one sentence, which appears in the novel.
An explanation
In which chapter does there exist a sentence in the novel <the novel> with the same or similar meaning as "<a sentence not from the original text>"? Please output the chapter name or index.
A chapter index or title
<A character> is used to be <positive or negative> and finally becomes a <negative or positive> one in the novel. Tell in one sentence which episode marks this character’s change.
A plot in one sentence
chara
Who are mentioned with names in the <an organization, a family, or a club> in the novel?
A list of names
Please list 3 aliases or designations of <a character> in the novel.
3 aliases or designations
Who is <a minor character, or a character without name, or a character that appears only once> in the novel?
A description of character
settg
In which <cities, or countries> does this story take place in the novel?
A list of cities or countries
In which year does the earliest event happen, and in which year does the latest even happen in the novel?
A range in years
relat
What is the relationship between <a character> and <an alias or a nickname of this character> in the novel?
A relationship
Table 7: Question template adopted in data annotation. Question types include character (chara), meaning (mean), plot, relation (relat), setting (settg), times and span. ’<>’ label indicates the entity for annotators to fill in.
The 17 pieces of question template adopted in the annotation procedure are presented in Table 7.
A.2 Distribution of Input Token Count
Figure 5: Token Count Distribution in NovelQA. The token count of both the novel and the questions are counted.
The input token count of each novel in NovelQA is calculated by adding the book length to the lengths of its related questions. The distribution of the token count is illustrated in Figure 5.
A.3 Data Classification
Dimension Type Percentage Description Example
Multi-hop 34.98%
Questions requiring knowledge across multiple paragraphs, or even multiple chapters, to be solved.
How many times has Bran jumped off and ran? in A Game of Thrones
By Complexity Single-hop 42.74%
Questions requiring knowledge from one or several adjacent single sentences to be solved.
According to the Colonel, what did he smoke in McQueen’s compartment? in Murder on the Orient Express
Detailed 22.19%
Questions requiring knowledge from one or two adjacent sentences to be solved. Detailed questions are distinguished from the single-hop class by involving information that is too minor to impact other plots, making the details difficult to be recalled.
How many candles Madame Magloire lighted when the Bishop had his last dinner with Jean Valjean? in Les Misérables
Times 20.07%
About the number of times that a character, location, or plot appears in the novel.
How many times has Kitty kissed Walter? in The Painted Veil
Meaning 15.86%
The understanding of certain sentences or metaphors, e.g., to interpret the relationship of a certain metaphor and the novel itself, or find a specific sentence according to a paraphrase provided by the annotator.
In which chapter does there exist a sentence with the same or similar meaning as ‘The Marquis responded, ‘You do me too much honor. In any case, I lean toward that assumption.”? in A Tale of Two Cities
By Aspect Span 1.47%
About the range of the novel setting. To be specific, they either ask about the starting and ending year of the story or require listing all the cities or countries that are involved in the story.
In which year does the earliest event happen, and in which year does the latest even happen? in Tess of the d’Urbervilles
Setting 11.44%
About the time or place settings, besides those in the span type, are classified in this type.
Where did Diana’s cousins leave for the Debating Club concert? in Anne of Green Gables
Relation 7.15%
About the relationship of multiple character entities. To be specific, they ask either about the relationship of a character and their alias or designation, or about the relationship between different characters.
What is the relationship between Jean Valjean and Ultime Fauchelevent? (designation) Who are members of ABC friends? (membership) in Les Misérables
Character 18.29%
About the information of characters, besides those in the relation type, are classified into this type.
Who is Miss Beirne? in Dubliners
Plot 25.62%
We define a plot as "some character does something for once". Questions that ask about the information of any plots are classified into this type.
What does Clarissa repair in preparation for the night’s party? in Mrs.Dallory
Table 8: Data Distribution
The data in NovelQA are classified into 3 complexity-levels and focus on 7 aspects. A detailed description of criteria and examples for each class is presented in Table 8.
A.4 Prompts
QA Setting Prompt
Generative You are a literature professor. I will provide you with the full text of a novel along with a series of questions. Please thoroughly analyze the novel ’s content to accurately respond to each of the following questions. Book title: <title>; Book Content: <content>; Book ends. Questions start here: N $\times$(Question: <question>); Questions end here. Try your best to answer the questions based on the given full text the novel. The answer should be in short with only one or several words. Your output format should be ’Answer0: <answer>Answer1: <answer>… Answern: <answer>’, each answer in one line without outputing the questions and other info.
Table 9: Prompts used in each setting of question answering.
We present all the prompts involved in either tests or evaluations in Table 9.
Appendix B Experiments
B.1 Inter-evaluator Agreement for GPT-4-as-Evaluator
We also perform an inter-evaluator agreement test between human-evaluators and the GPT-4-evaluator. The results of Cohen’s Kappa score are presented in Table 6.
B.2 Relation between Multi-choice Accuracy and Evidence Positions
Figure 6: Analysis of Accuracy in Multichoice Setting by Token Indexes: This figure illustrates the accuracy in the generative setting of NovelQA, plotted against the token indexes of relevant evidences in the novels. Each subplot represents a different model. The x-axis, reflecting token indexes, is folded on the right due to the long-tail distribution in the lengths of the selected novels.
Figure 7: Analysis of Accuracy in Multichoice Setting by Relative Token Positions: This figure illustrates the accuracy in the multichoice setting of NovelQA, plotted against the percentage position of each question’s relevant evidence within the novel. Each subplot represents a different model.
chara mean plot relat settg span times avg
dtl 69.39 30.10 85.84 53.12 80.95 - - 57.62
mh 76.81 88.24 87.50 79.83 91.67 52.94 45.79 60.22
sh 86.27 88.10 92.03 57.14 87.01 - - 88.64
avg 80.81 55.46 90.36 72.73 85.98 52.94 45.79 71.80
(b) Claude 2.1
chara mean plot relat settg span times avg
mh 71.88 76.47 85.00 70.69 78.26 51.52 47.71 58.17
sh 82.28 76.72 83.29 50.00 80.89 - - 81.19
dtl 63.51 38.12 76.92 68.75 79.03 - - 58.02
avg 76.80 54.55 82.21 68.52 80.17 51.52 47.71 66.78
(c) InternLM2-Chat-7b
chara mean plot relat settg span times avg
mh 42.19 38.24 45.90 46.15 65.22 39.39 42.92 43.87
sh 44.44 39.34 44.56 28.57 48.15 - - 44.23
dtl 52.63 26.24 55.14 31.25 59.68 - - 41.54
avg 45.69 31.84 46.79 41.72 52.63 39.39 42.92 43.51
(a) GPT-4
chara mean plot relat settg span times avg
mh 55.88 44.12 53.12 56.78 66.67 44.12 38.53 45.15
sh 53.94 53.97 58.94 35.71 62.15 - - 57.26
dtl 47.96 25.98 55.75 18.75 55.56 - - 30.00
avg 52.86 37.36 57.70 47.56 60.98 44.12 38.53 49.18
(d) InternLM2-Chat-20b
Besides the evaluation in generative settings, we also prompted the models to collect their responses for the multichoice versioned questions. Table 10 presents the accuracies of four models in multichoice settings in each question type. Figure 6 and Figure 7, presenting the relationships between the accuracy and the absolute or relative positions accordingly, show similar trends to which are observed in the generative setting.
Figure 8: Performance Analysis by Relative Positions within Different Token Ranges: This figure illustrates the accuracy of various long-context LLMs in the multichoice setting, segmented by novels’ length categories: 65k-100k tokens and over 100k tokens. It highlights the models’ performance trends relative to the evidence position within these ranges, showcasing the ‘lost-in-middle’ phenomenon in shorter texts and a performance decline towards the end in longer texts, thus revealing distinct comprehension challenges faced by LLMs when processing texts of varying lengths.
B.3 Representative Errors
Fault Type Subtype Setting Example QA
Overlooking refers to the model’s neglect of details. As mentioned above, the questions in detailed category involve minor characters, plots, or settings. Diving further, the reasons why these details are difficult to be recalled lie in two aspects: (1) They do not contribute to the character development, other plots, or the main themes, and thus reading the rest of the novel does not help to remind this detail; (2) Since most novels have derivative works (e.g., films, fan works, and book reviews), where the detailed information is eliminated to form a condensed narration. As the derivative works spread further and appear more frequently in the model’s training data, they have a higher probability of becoming the models’ inner knowledge, which is similar to (Chang et al., 2023a)’s observation, and vice versa for those omitted details. These two factors contribute to the difficulty in the model’s recalling details and thus result in overlooking errors.
Miscounting Researches (Li et al., 2023a; Feng et al., 2023; Wang et al., 2021b) has revealed shortcomings in the counting ability of LLMs, especially autoregressive-decoder-based models, and methods unfolding the outputs such as chain-of-thought prompting can enhance their counting ability. Our test does show that models make mistakes with numbers. Though the errors in the case of generative responses may be due to not following instructions and simply outputting "multiple times" instead of the desired specific times, the accuracy in the multichoice setting has still only reached 38.53% to 49.56% for the chosen four models, as shown in Table 4. Even in the simplest question which asks for the appearing frequency of certain phrases, the models still make mistakes.

### v2

@@ D.3 Data Classification
Table 7: Data Distribution of NovelQA. Each question is labeled with a Complexity and an Aspect label. The Complexity dimension comprises three categories, multi-hop, single-hop, and detail, while the Aspect dimension comprises seven categories including times, meaning, and so forth.
Dimension Type Percentage Description Example
Multi-hop 34.98% Questions requiring knowledge across multiple paragraphs, or even multiple chapters, to be solved. Has Crusoe been on a voyage? If so, how many times has this plot happened? in The Life and Adventures of Robinson Crusoe
By Complexity Single-hop 42.74% Questions requiring knowledge from one or several adjacent single sentences to be solved. According to the Colonel, what did he smoke in McQueen’s compartment? in Murder on the Orient Express
Detail 22.19% Questions requiring knowledge from one or two adjacent sentences to be solved. Detail questions are distinguished from the single-hop class by involving information that is too minor to impact other plots, making the details difficult to recall. How many candles Madame Magloire lighted when the Bishop had his last dinner with Jean Valjean? in Les Misérables
Times 20.07% About the number of times that a character, location, or plot appears in the novel. How many times has Kitty kissed Walter? in The Painted Veil
Meaning 15.86% The understanding of certain sentences or metaphors, e.g., to interpret the relationship of a certain metaphor and the novel itself, or find a specific sentence according to a paraphrase provided by the annotator. In which chapter does there exist a sentence with the same or similar meaning as ‘The Marquis responded, ‘You do me too much honor. In any case, I lean toward that assumption.”? in A Tale of Two Cities
By Aspect Span 1.47% About the range of the novel setting. To be specific, they either ask about the starting and ending year of the story or require listing all the cities or countries that are involved in the story. In which year does the earliest event happen, and in which year does the latest even happen? in Tess of the d’Urbervilles
Setting 11.44% About the time or place settings, besides those in the span type, are classified in this type. Where did Diana’s cousins leave for the Debating Club concert? in Anne of Green Gables
Relation 7.15% About the relationship of multiple character entities. To be specific, they ask either about the relationship of a character and their alias or designation, or about the relationship between different characters. What is the relationship between Jean Valjean and Ultime Fauchelevent? (designation) Who are members of ABC friends? (membership) in Les Misérables
Character 18.29% About the information of characters, besides those in the relation type, are classified into this type. Who is Miss Beirne? in Dubliners
Plot 25.62% We define a plot as "some character does something for once". Questions that ask about the information of any plots are classified into this type. What does Clarissa repair in preparation for the night’s party? in Mrs.Dallory
Question Type Abilities Required
Multi-hop Ability to retrieve and integrate scattered information.
Single-hop Ability to retrieve information and summarize.
Detail Ability to precisely identify specific, subtle details.
Times Ability to retrieve facts, integrate facts, and reason.
Meaning Ability to retrieve ambiguous information and reason.
Span, Setting, Relation Ability to retrieve facts, integrate facts, and reason.
Character, Plot Ability to subtle or ambiguous information and reason.
Table 8: Question types and their corresponding LLM abilities. Each type of question in NovelQA is designed to test one or more abilities of LLMs, including the abilities to retrieve information, to identify details, and so forth.
The data in NovelQA are classified into 3 complexity levels and focus on 7 aspects. Table 7 presents a detailed description of each class’s criteria and examples.
D.4 Inter-Annotator Agreement Tests
We have conducted two separate Inter-Annotator Agreement (IAA) tests in this work.
The first IAA experiment focuses on evaluating the quality of annotated question-answer pairs. Annotators are required to choose books they are familiar with but have not annotated themselves to answer questions on. As annotators are mostly from the same local university and share similar courses and projects, many can find books they have read in common. We select books read by at least two annotators and have the other reader answer multichoice questions. As the respondents are quite familiar with the target novels and have a strong academic background in English Literature, it takes them around 2-3 hours or less to complete each novel.
The second IAA test assesses model answer evaluation by calculating the correlation between human evaluators and GPT-4 evaluators. Annotators also serve as human evaluators for model-generated answers to the novels they were familiar with. Their evaluations were compared to those of GPT-4 evaluators, with Cohen’s kappa score calculated to measure agreement. Human evaluators, familiar with the content, typically completed their reviews in under half an hour per novel. We also perform an inter-evaluator agreement test between human-evaluators and the GPT-4-evaluator. The results of Cohen’s Kappa score are presented in Table 4.
D.5 Prompts
Table 9: Prompts used in each setting of question answering. In each prompt, the model is assigned the identity of a literature professor reviewing student answers, with a detailed task description and clear specifications for the input and desired output. The angle brackets ‘<>’ indicate the contents varying among each input.
QA Setting Prompt
Generative You are a literature professor. I will provide you with the full text of a novel along with a series of questions. Please thoroughly analyze the novel’s content to accurately respond to each of the following questions. Book title: <title>; Book Content: <content>; Book ends. Questions start here: N $\times$(Question: <question>); Questions end here. Try your best to answer the questions based on the given novel full text. The answer should be in short with only one or several words. Your output format should be ’Answer0: <answer>Answer1: <answer>… Answern: <answer>’, each answer in one line without outputing the questions and other info.
We present all the prompts involved in both test and evaluation phases in Table 9. The prompts generally follow the formula of zero-shot prompting. In each prompt, the model is firstly assigned the identity of a literature professor reviewing student answers. Then the prompt provides the model with a detailed task description and clear specifications for the input and desired output.
Appendix E Experiments and Analysis
Table 10: Model Performance Analysis Pre and Post 100K Tokens: The accuracies of questions on evidences pre-100K and post-100K are calculated separately and compared, where the ‘100K’ indicates the aboslute starting token position of evidences relating to answer. The result shows an obvious contrast as the accuracy drops sharply after the 100K, in both generative and multichoice settings on each model.
Multichoice Generative
pre-100K post-100K pre-100K post-100K
GPT-4 83.17 59.22 53.64 32.01
Claude2.1 77.65 54.46 61.96 30.57
InternLM-7b 51.78 38.17 37.71 18.23
InternLM-20b 55.59 40.87 37.74 21.10
Weighted Avg. 66.24 48.08 46.84 25.36
(a) GPT-4 (b) Claude 2.1
chara mean plot relat settg span times avg chara mean plot relat settg span times avg
mh 55.88 44.12 53.12 56.78 66.67 44.12 38.53 45.15 mh 71.88 76.47 85.00 70.69 78.26 51.52 47.71 58.17
sh 53.94 53.97 58.94 35.71 62.15 - - 57.26 sh 82.28 76.72 83.29 50.00 80.89 - - 81.19
dtl 47.96 25.98 55.75 18.75 55.56 - - 30.00 dtl 63.51 38.12 76.92 68.75 79.03 - - 58.02
avg 52.86 37.36 57.70 47.56 60.98 44.12 38.53 49.18 avg 76.80 54.55 82.21 68.52 80.17 51.52 47.71 66.78
(c) InternLM2-Chat-7b (d) InternLM2-Chat-20b
chara mean plot relat settg span times avg chara mean plot relat settg span times avg
mh 42.19 38.24 45.90 46.15 65.22 39.39 42.92 43.87 mh 76.81 88.24 87.50 79.83 91.67 52.94 45.79 60.22
sh 44.44 39.34 44.56 28.57 48.15 - - 44.23 sh 86.27 88.10 92.03 57.14 87.01 - - 88.64
dtl 52.63 26.24 55.14 31.25 59.68 - - 41.54 dtl 69.39 30.10 85.84 53.12 80.95 - - 57.62
avg 45.69 31.84 46.79 41.72 52.63 39.39 42.92 43.51 avg 80.81 55.46 90.36 72.73 85.98 52.94 45.79 71.80
E.1 Implementations
We also tested Gemini-1.5555https://deepmind.google/technologies/gemini/flash/, but generation for 816 questions from 29 novels was blocked for unknown reasons (with only 14 of these novels under copyright protection). Consequently, we have decided not to present Gemini’s result.
E.2 Multi-choice Performance on Each Type of question
Besides the evaluation in generative settings, we also prompted the models to collect their responses for the multi-choice versioned questions. Table 11 presents the accuracies of four models in multi-choice settings in each question type.
E.3 Relation between Multi-choice Accuracy and Evidence Positions
@@
Figure 6: Analysis of Accuracy in multi-choice Setting by Absolute and Relative Token Positions: This figure illustrates the accuracy in the multi-choice setting of NovelQA, plotted against the absolute position (left) and percentage position (right) of each question’s relevant evidence within the novel. Each subplot represents a different model. The x-axis of the absolute position figure (left), reflecting token indexes, is folded on the right due to the long-tail distribution in the lengths of the selected novels.
Figure 6 which presents the relationships between the accuracy and the absolute or relative positions accordingly shows similar trends to which are observed in the generative setting. To be specific, the accuracy by absolute token position remains high when related evidences are before 100K’s text length, while drops after 100K. Meanwhile, the accuracy by relative position remains relatively even. A comparison is made between the accuracies within two ranges, 65K(the lowest token count) to 100K (namely pre-100K) and 100K to the end (namely post-100K). Figure 7 and Table 10 present a clearer contrast between these two ranges, where the accuracies drop sharply after the 100K token.
Figure 7: Performance Analysis by Relative Positions within Different Token Ranges: This figure illustrates the accuracy of various long-context LLMs in the multi-choice setting, segmented by novels’ length categories: 65K(the lowest token count) to 100K tokens and over 100K tokens. It highlights the models’ performance trends relative to the evidence position within these ranges, showcasing the ‘lost-in-middle’ phenomenon in shorter texts and a performance decline towards the end in longer texts, thus revealing distinct comprehension challenges faced by LLMs when processing texts of varying lengths.
E.4 Representative Errors
Fault Type Subtype Setting Example QA
Overlooking refers to the model’s neglect of details. As mentioned above, the questions in detailed category involve minor characters, plots, or settings. Diving further, the reasons why these details are difficult to be recalled lie in two aspects: (1) They do not contribute to the character development, other plots, or the main themes, and thus reading the rest of the novel does not help to remind this detail; (2) Since most novels have derivative works (e.g., films, fan works, and book reviews), where the detailed information is eliminated to form a condensed narration. As the derivative works spread further and appear more frequently in the model’s training data, they have a higher probability of becoming the models’ inner knowledge, which is similar to [Chang et al., 2023]’s observation, and vice versa for those omitted details. These two factors contribute to the difficulty in the model’s recalling details and thus result in overlooking errors.
Miscounting Researches [Li et al., 2023a, Feng et al., 2023] has revealed shortcomings in the counting ability of LLMs, especially autoregressive-decoder-based models, and methods unfolding the outputs such as chain-of-thought prompting can enhance their counting ability. Our test does show that models make mistakes with numbers. Though the errors in the case of generative responses may be due to not following instructions and simply outputting ‘multiple times’ instead of the desired specific times, the accuracy in the multi-choice setting has still only reached 38.53% to 49.56% for the chosen four models, as shown in Table . Even in the simplest question which asks for the appearing frequency of certain phrases, the models still make mistakes.
@@ E.4 Representative Errors
E.5 Close-book Question Answering
@@
We employ a Close-book QA scenario to assess the extent to which models rely on the content of novels versus using their internal knowledge to answer questions. In this approach, the models are not given access to the text of the novels and must rely solely on their pre-existing knowledge to provide answers. Given that our selected novels are well-known and representative of their genres, it’s inevitable that LLMs have encountered their texts during training and retained some of their content. The results of this evaluation are presented in Table 13.
@@
Models like GPT-4 and Claude 2.1 achieve notable scores in the Close-book setting (60.94% and 51.77% in multi-choice, 34.30% and 22.36% in generative, respectively) indicating that they have internalized significant portions of the novels’ content during training. This internal knowledge allows them to perform reasonably well even without direct access to the text. The difference in performance between the Close-book and standard settings underscores the challenges in long-context understanding. While models can retain and recall information from well-known texts, their ability to comprehend and use such information to answer questions accurately diminishes in the absence of the text. This suggests that long-context understanding, as measured by the main results, might still be more challenging than it appears, as models benefit from having the text directly available.